OpenStudy (anonymous):

A stone dropped into a still pond sends out a circular ripple whose radius increases at a constant rate of 2.8 ft/s. (a) How rapidly is the area enclosed by the ripple increasing when the radius is 4 feet?

OpenStudy (anonymous):



OpenStudy (anonymous):

It says the radius is 4 ft. the rate it is increasing is 2.8 ft/s

OpenStudy (anonymous):

so it's asking us how fast is the area being enclosed. so we will need to know the equation for area of a circle

OpenStudy (anonymous):

oh. A=pir^2. Am I suppose to take the derivative of the equation so 2pir(dr/dt)?

OpenStudy (anonymous):

Yeeeeah lol

OpenStudy (anonymous):

Related Rates

OpenStudy (anonymous):

I'm bad at this stuff. repost this and get some help from another person.

OpenStudy (anonymous):

it's okay. lol Neither am I. :)

OpenStudy (anonymous):

Thanks, though!

OpenStudy (ranga):

A = (pi)(r^2) dA/dt = 2(pi)rdr/dt Put r = 4 and dr/dt = 2.8 and find dA/dt.

OpenStudy (anonymous):

Thanks! I got that!!

OpenStudy (ranga):

cool. yw.

OpenStudy (anonymous):

uh if it asks for the area after 7.2 seconds, where would I put it?

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

*put the time?

OpenStudy (ranga):

dr/dt = 2.8 Integrate both sides and express r as a function of time t first.

OpenStudy (ranga):

dr/dt = 2.8 Integrate both sides: r = 2.8t + C when t = 0, r = 0 0 = 0 + C. So C = 0 r = 2.8t At time t = 7.2 find r. Then Area = pi(r^2)

OpenStudy (anonymous):

Ohh. Okay. Yes, I get it! Thanks!!

OpenStudy (ranga):

you are welcome.